OpenStudy (aaronandyson):

Use mathematical induction to prove De Moivre's theorem [ R (cos t + i sin t) ] n = R n(cos nt + i sin nt) for n a positive integer.

OpenStudy (aaronandyson):

@Preetha @pooja195 @sarahnichols3 @BAdhi @iambatman

OpenStudy (aaronandyson):

@Marki

OpenStudy (anonymous):

`this is the theorem ` \([ R (\cos ~t + i ~\sin ~t) ]^n = R^n(\cos~ nt + i~ \sin~ nt) \) make sure next time to type it fully

OpenStudy (anonymous):

maybe @mathmath333 free to solve it for u

OpenStudy (anonymous):

note that \([ R (\cos ~t + i ~\sin ~t) ]^n = R^n(\cos~ nt + i~ \sin~ nt)=e^{nR\theta}\)

OpenStudy (aaronandyson):

[ R (cos t + i sin t) ] k R (cos t + i sin t) = R k(cos kt + i sin kt) R (cos t + i sin t) ?

OpenStudy (aaronandyson):

(cos kt cos t - sin kt sin t) = cos(kt + t) = cos(k + 1)t (sin kt cos t + cos kt sin t) = sin(kt + t) = sin(k + 1)t ????

OpenStudy (badhi):

\[(\cos kt + i \sin kt )(\cos t + i \sin t) =\\ [ \cos (kt)\cos t - \sin(kt)\sin t]+i[\sin(kt)\cos t+ \cos(kt)\sin t]\] Use above identity of yours

OpenStudy (aaronandyson):

[ R (cos t + i sin t) ] k + 1 = R k + 1 [ cos (k + 1)t + sin(k + 1)t ]

OpenStudy (badhi):

Please use brackets where needed tell us what else you need to know and where you are stuck :)

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{ (\cos x + i \sin x)^n = (\cos nx + i \sin nx) \hspace{.33em}\\~\\ \normalsize \text{for n=1} \hspace{.33em}\\~\\ (\cos x + i \sin x)^n \hspace{.33em}\\~\\ = (\cos x + i \sin x)^1 \hspace{.33em}\\~\\ = (\cos x + i \sin x)\hspace{.33em}\\~\\ \normalsize \text{True} \hspace{.33em}\\~\\ \normalsize \text{assume it is true for n=k, so} \hspace{.33em}\\~\\ (\cos x + i \sin x)^k = (\cos kx + i \sin kx)------(1) \hspace{.33em}\\~\\ \normalsize \text{prove that it is also true for (n=k+1) such that } \hspace{.33em}\\~\\ (\cos x + i \sin x)^{k+1} = (\cos (k+1)x + i \sin (k+1)x) \hspace{.33em}\\~\\~\\ \underline {~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ }\hspace{.33em}\\~\\ (\cos x + i \sin x)^{k+1} \hspace{.33em}\\~\\ =(\cos x + i \sin x)^{k}(\cos x + i \sin x) \hspace{.33em}\\~\\ =(\cos x + i \sin x)^{k}(\cos x + i \sin x) \hspace{.33em}\\~\\ =(\cos kx + i \sin kx)(\cos x + i \sin x) -------(from~~1)\hspace{.33em}\\~\\ =\cos kx\cdot \cos x + \cos kx\cdot i\cdot \sin x +i\cdot \sin kx\cdot \cos x+i\cdot \sin kx\cdot i \sin x\hspace{.33em}\\~\\ =\cos kx\cdot \cos x -\cdot \sin kx\cdot \sin x + i(\cos kx\cdot \cdot \sin x +\cdot \sin kx\cdot \cos x)\hspace{.33em}\\~\\ =\cos (kx+x) + i(\sin kx+x)\hspace{.33em}\\~\\ =\cos x(k+1) + i(\sin x(k+1))\hspace{.33em}\\~\\ }\end{align}\)